The membership fee for joining a garden association is $24 per year local botanical garden charges members of the gardening Association three dollars for admission to the garden.Nonmembers of the association are charged six dollars. After how many visits to the garden is a total cost members including the membership fee the total cost for nonmembers
Answer by rfer(16322)

24+3x=6x
24=6x-3x
24=3x
24/3=x
8=x
----------
24+3*8=6*8
48=48